### v1

@@ Appendix E Machine learning with the quantum reservoir
We then computed the mean squared error between the resulting $y_{\text{prediction}}$ and label for the training point that produced the underlying reservoir feature, and used back-propagation to compute the gradient for our linear layer. The linear layer was then updated using the ADAM optimizer [ADAMpaper] with the default settings of $\beta_{1}=0.9,\beta_{2}=0.999$ and a learning rate of $0.01$. For our reservoirs, we tried both methods of training the linear layer and used whichever yielded the best accuracies. Empirically, we found that while pseudo-inverse training was better in some cases, training the linear layer with back-propagation often yielded quite large accuracy advantages over pseudo-inverse.
Appendix F Supplementary information machine learning tasks
@@
In this section, we discuss about the algorithm for generating the dataset for the classification of digital modulation schemes on radio signals. The digital modulation scheme involves encoding sequences of binary values into the amplitude and phase of a radio signal for a fixed duration. The number of binary values encoded depends on the modulation scheme. For example, for BPSK (binary phase shift key), each symbol (change in property of the signal) encodes one bit of information. For $32$QAM (quadrature amplitude key), there are $32$ possible values, which allows each symbol to contain $5$ bits of information. For this task, we keep the symbol rate fixed across all the tasks. Moreover, the pulses generated by the arbitrary waveform generator (AWG) all occur at the baseband frequency. This signal is then upconverted to the frequency of the cavity before sent to the device. To generate the set of possible sequences, we randomly select each symbol with equal probability. This corresponds to the case of each possible binary string of digits encoded to be equally likely. Due to memory constraints on the AWG, we cannot output a continuous encoded signal for long durations, corresponding to the regime of large samples of the reservoir. We circumvent this constraint by realizing that, for this task, there are no correlations in the encoded binary digit sequence (since each symbol is equally likely). Therefore, for example, the probability of a long binary digit sequence can be correctly emulated by sampling multiple short binary digit sequences and concatenating them together. For this task, we can simply achieve this by generating a signal with eight symbols, which is the number of symbols enter our QRC before its state is completely reset.
@@
Figure M14: Histogram of integrated value of the input signal for the noise classification task. In this work, we enforced a normalization condition on the amplitude of the filter functions. We set the normalization such that the long time integral of the signal corresponds to a value with zero mean and the same standard deviation. This is visually seen from the probability density function from the dataset of the classes of noisy signals. We do this to ensure that any reservoir which simply integrated the signal, before applying a non linear kernel, cannot classify the different signals. The fact that our reservoir is able to solve this task can therefore be associated with the continuous-time processing by the cross-Kerr interaction between the oscillator the cavity. Enforcing this normalization is mathematically equivalent to setting the DC component of the filter functions to be the same value in frequency space.
To generate the dataset describing the task of classifying noisy signals using the QRC (see Fig. 4), we start with emulating white noise. At each time step of the sampling rate of the AWG, we choose a value for in-phase and quadrature signals uniformly between the unit interval (up to an overall normalization). While this is limited to the sampling rate of the AWG (around $2\times 10^{9}$ samples per second), this is much larger than any relevant time scale of the experiment. Therefore the approximation of broadband white noise is appropriate to describe the effect of the signal on the system. We then apply “kernels” as a convolution in the time domain to each need seed of the white noise generated signal. This can also be thought of a bandpass filtering function in frequency domain. The classification task is then to identify the kernel. Each kernel is defined by a time domain function. The only hyper-parameter to describe each kernel is the overall scaling value. In this work, we set the DC component of this kernel in frequency domain to be the same for all classes (set to unit value without loss of generality). In the time domain, this corresponds to scaling the amplitude such that the area enclosed by the filter function in time is the same for all functions.. We do this to make sure that a direct integration of the signal over a time domain much longer than the correlation length introduced by the kernel, cannot distinguish the signals from each (see Fig. M14). The above normalization ensures the random variable associated with the value of this integrated value is the same for all distributions. Therefore, this ensures that any ability of the reservoir to classify the signals is intrinsically from its computational capacity to distinguish short-time correlations (in this work we choose a correlation time scale of $50$ns and $600$ns, with kernel functions of Gaussian, Lorentzian, and the Inverse function: generating a total of $6$ classes.).
Appendix G Simulation of the quantum reservoir
@@ Introduction
Figure M15: Classification accuracies obtained from numerical simulations of the QRC. Numerical simulations of the quantum reservoir can help guide the expected performance of the system in experiment. Here, we simulate the three main tasks considered in the paper: Spiral (a time independent input with the goal of classifying the arms of the spiral), Digital Modulation (a slow varying time dependent signal with the goal of identifying the digital modulation scheme used in the signal), Correlated Noise (a fast varying time dependent signal with the goal of classifying the “kernel” of correlation function).
@@
To make the simulations efficient, we make certain assumptions on the quantum system. Firstly, we treat the reservoir controls of qubit rotations and conditional displacements with a “gate”-based unitary. However, to take into account the analog, continuous dynamical evolution implemented by the cross-Kerr interaction term in the Hamiltonian, the interval of the input into the system is implemeted with the full time-dependent Hamiltonian evolution (using QuTiP’s “mesolve” functions). Finally, another approximation we make (in favor for simulation speed) is ignoring decoherence effects. To ensure this approximation is valid, we performed simulation with the stochastic wavefunction approach with photon loss and qubit dephasing rates measured in experiments [PhysRevLett.68.580]. We obtain differences in expected classification accuracies within error bars (which are obtained from different datasets from repeated simulations). This also gives us confidence that the role of decoherence in the system plays a minimal role in the computational capacity of the reservoir. The results of the simulations with the third central moment are plotted in Fig. M15. Interestingly, the performance as a function of the number of samples agrees to experiment within the same order of magnitude. This gives a good estimate for the experimental time required for produce a classification accuracy versus shots curves in experiments. For all the three tasks, the reservoir approaches $100\%$ with sufficient samples or integration time of the input.
@@
Figure M16: Simulations comparing the performance of a continuous-time quantum reservoir (our work), and a discrete-time qubit-based quantum reservoir, based on a recently implemented protocol [7]. (a.) Schematic reservoir protocol for our experiment (on the left), versus that introduced in  [7], involving two qubits as the data qubits, and two as the ancillas. The reservoir consists of single qubit rotations interleaved with CNOT gates between the data qubits and data and ancilla qubits. In this simulation, we simulate the performance of the two reservoirs as a function of the delay $\tau$ between two durations of the input. Such a delay can be introduced by the finite pulse durations and latencies introduced by the FPGA. (b.) Classification accuracy cueves for the two reservoirs as a function of shots, for different values of delays. The continuous time analog reservoir is much more robust to delays between inputs compared to the discrete time reservoir implementation. (c.) Plot of accuracies at $5000$ shots for the two reservoir implementations as a function of delay time. Experimentally relevant times include $4\mu s$ for our experiment, and around $8\mu s$ [hua2023exploiting] for the experimental realization of [7] on an IBM quantum computer.
@@
In this section, we benchmark the performance of our continuous-time-continuous variable QRC in comparison with other hardware implementations of reservoirs. To highlight the benefit of our QRC in processing time varying input signal, we compare the simulation of our reservoir with that of a recent QRC scheme involving repeated measurements on a multi-qubit based superconducting circuit quantum system [7]. For this comparison, we simulate the expected performance of both systems on the task of classifying different noise signals with the classes described in Fig. 4. While our reservoir can naturally interface with analog signal, this is not the case with the protocol introduced in [7]. For this simulation, the signal is samples at discrete time and input to the system as a scalar parameter (one for the in-phase value and one for quadrature value). To highlight the advantage of our QRC, we slightly modify the task introduced in the Fig. 4. Here, we normalize the six filter functions such that the integral of the filter function in frequency domain is kept constant. We do this such that the standard deviation associated with the distribution of the sampled signal is the same across all signals. The only information distinguishing amongst the signals is in the correlation between two close samples in time. To elucidate this reasoning, we simulation the performance of the two reservoirs as a function of the time duration in between two samples of the signal (in the case of the discrete qubit based reservoir) and integration windows (for our analog reservoir). Such a finite duration can arise from finite-pulse durations of reservoir protocols, qubit-measurement times, and the finite latency introduced by the classical FPGA processor. For example, for our experiment, this time is around $4\mu s$, mostly arising from the measurement of the qubit and the parity of the cavity. In experiment (Fig. 4), we had generated and timed the input wave forms such that the delay between inputs is essentially $0\mu s$. For a typical IBM quantum device with mid-circuit measurement, the protocol used in Ref. [7], the finite latency can be estimated to be around $8\mu s$ [hua2023exploiting]. The protocol for the discrete-time quantum reservoir is designed to only act on real values input signals. However, for a continuous signal in the rotating frame, we have both the in-phase and quadrature values. For the experimental quadrature, these values correspond to displacements on the oscillator in orthogonal directions. To extend the scheme presented in [7], we do the following minimal change: we interleave the between sample points of in-phase and quadrature values. We could have chosen these points with the delay of $\tau$ in between each. However, this might have the effect of introducing twice the delay compared to the continuous time reservoir. Therefore, we choose the relaxed constraint of the input such that both the in-phase and quadrature values are chosen at the same point, with just a delay in between two different in-phase and quadrature points.
@@
Figure M17: Simulations of the performance of the cavity-and-qubit reservoir system with its components: just a cavity reservoir and just a qubit reservoir. (a.) The circuit diagrams of the three quantum reservoir simulated here. 1) The circuit diagram of the experimental QRC in this work: a cavity coupled to a qubit. The input is interfaced to the cavity, and the state of the qubit is measured using the standard dispersive readout technique with a readout resonator. 2) The circuit diagram of a cavity as a reservoir. In this case, the natural output of the cavity is the transmitted signal. 3) The circuit diagram of a qubit reservoir. In this case the input is coupled instead to the qubit. (b.) Simulation of the classification accuracy of the three different reservoirs for the three tasks considered in the paper. As a fair comparison, the output feature vector dimension is kept constant for all three reservoir. The experimental setup drastically outperforms either of its components, just a cavity (in (c.) and just a qubit (in (d.)), which highlights the important role of entanglement in classification accuracy.
@@
A cavity coupled to a qubit is a hardware efficient quantum system to perform reservoir computing on analog signals. In this section, we motivate this by simulating the performance of other natural choices of quantum reservoirs: a single qubit, and a single cavity. The protocol for these systems are inspired by what one can naturally perform in experiment. To make a reservoir with a cavity, we couple the input into the cavity (as is the case for the experimental design). To readout the cavity, we perform a transmission style Homodyne measurement, which infers the mean field value of the cavity. This is a continuous form of measurement, where the output feature is a time dependent radio frequency signal at the frequency of the cavity mode. Since the cavity is always in a coherent state, the output time trace is linearly dependent on the incoming signal. For a fair comparison, we only use a handful of values from the time trace (as many as the number of measurements in the experiment). While this might seem restrictive, we process this via the same method as the case of the experimental reservoir, by computing the functional definition of the central moments. This does not necessarily make sense for this protocol, since the output do not correspond to samples from a discrete probability distribution, but can nevertheless introduce non-linearities in the representation of the feature vector. These non-linearities can improve the performance of the reservoir beyond a linear layer. This is observed for the case of time-indepedent Spiral classification, where the cavity reservoir performs better than random. This performance is solely due to the “post-processing” of the output of the reservoir we adopt for our experiment. However, for time dependent tasks, the performance is hardly better than random.
@@
Another natural candidate is a single qubit reservoir. For this case, we directly interface the signal to the qubit. A qubit is able to naturally represent non-linear functions of the input, which can be intuitively seen by visualizing the action of a qubit rotation on a Bloch sphere. As a fair comparison, we choose the same qubit reservoir controls in experiment, which involve qubit pulses before, during and after the continuous input. The output is a string of binary outcomes of qubit measurements, which can be done experimentally with the use of a readout resonator. Each reservoir of the qubit lasts twice as long as the experimental QRC to obtain the same feature vector size. This is then processed the same as the cavity qubit coupled reservoir, before applying a trained linear layer. Interestingly, the qubit fails to perform better than random for the Spiral task. On the other hand, it is able to reach near $100\%$ accuracy for the time dependent signal classification tasks. The ability for even a single qubit to successfully perform a many-class classification task is illuminating at the remarkable processing capabilities of reservoir. However the total input signal required can be more than order of magnitude longer compared to that for the experimental QRC to achieve the same accuracy. The ability of a cavity coupled to a qubit system to perform significantly better than either of its components provides a clear picture of the important role entanglement can play.
@@
Figure M18: Exploring the computational capacity of more complex reservoirs, involving a single cavity mode coupled to multiple qubits. (a.) The reservoir protocol, before measurement. The protocol is inspired by naturally extending the protocol of the experiment. The state of each qubit, along with the parity of the cavity, is sampled afterwards. (b.) Schematic illustration of the state of the cavity after the generalized qubits conditioned cavity displacement implemented in this reservoir, for the case of two qubits. Here the cavity is displaced by a unique value of each $2^{N}$ combinations of qubit possibilities, for a reservoir with N qubits. (c.) To illustrate the computational capacity of such reservoir, we simulate the quantum system to obtain the classification accuracy as a function of the duration of signal received for the task of classifying correlated noise signals. Increasing the number of qubits generally increases the performance of the reservoir. For different values of shots of the reservoir, which corresponds a total time of input, we plot the classification error as a function of qubits. The case of zero qubits in the reservoir corresponds to just a cavity reservoir.
@@
Quantum reservoir computing is a promising paradigm in the NISQ era. It is therefore interesting to consider the potential benefits in performance with larger devices, which are within reach of today’s experimental capabilities. As a natural extension of our quantum reservoir, we consider a scenario of one continuous variable cavity mode dispersively coupled to multiple qubits. For simplicity, we assume the dispersive strength of each qubit to the cavity is the same. To motivate the capacity of such a reservoir, we simulate the system for upto four qubits to estimate the classification accuracy for the task of identifying correlated noise signals. The unitary protocol is illustrated in Fig M18 (a.). The protocol begins with each a $\pi/2$ pulse on each qubit, which brings the state of the qubit onto the equator of the Bloch sphere. To entangle the qubits with cavity, we simulate the action of a generalized qubits conditioned cavity displacement. This involves a displacement on the cavity, whose value is different for each $2^{N}$ N-qubit possibilities. An example of the action of this operator is depicted in Fig M18 (b.) for the case of two qubits. Here there are four possibilities of the states of the qubit, and each is associated with a displacement value of the four corners of the square. The choice of the displacement values was chosen somewhat arbitrarily, but serves to highlight a situation of an efficient multi-component entanglement. For the case of two qubits, the real and imaginary components of the displacement where either ±$\pm 0.5$. For the case of three qubits, there are eight total possible states. The set of displacements chosen form a three by three grid state, ranging between ±$\pm 1$, excluding the center of this grid (which is centered at the origin). For the case of four qubits, a four by four grid uniformly distributed between ±$\pm 1.5$ cover all sixteen possibilities. The choice of which displacement corresponds to which qubits state was also made somewhat arbitrary (the motivation is even without much design choices, a reservoir can successfully implement machine learning!). For these simulations, each position of the grid is associated with a decimal value, increasing sequentially from left to right, starting from the top left and progressing towards the bottom right (starting with zero). The associate qubits state the displacement is conditioned on is the binary string representation of the decimal value, with a sequence length of the number of qubits.
@@
Fig M18 (c.) is the classification accuracy as a function of the total time of input signal received for the reservoirs with different number of qubits, for the task of noise classification. While they all achieve essentially $100\%$ accuracy, the total time required to achieve this accuracy drops significantly with increasing number of qubits. Other than the single qubit reservoir (the experimental protocol), the performance of the reservoir is similar both at the low and high signal duration regime, differing only in the intermediate regime. The reason for the difference in behavior of the performance for the case of cavity coupled to a single qubit is the slight change in reservoir protocol. To accurately account for the experimental protocol, the state of the qubit is determined by the outcome of the parity measurement of the cavity. This was not implemented in the simulations for multiple qubits. This ends up improving the performance of this reservoir for this task in the low signal duration regime. However, in the higher signal duration regime, increasing the number of qubits increases the accuracy.
@@
The classification error as a function of number of qubits in the reservoir is plotted in Fig M18 (d.), including the case for just a cavity (zero number of qubits in the reservoir), for a select number of total shots of the entire reservoir. Very crudely, the error in classification seems to reduce exponentially with every additional qubit in the reservoir.
Appendix H Theoretical analysis of the expressivity of our QRC for time-independent signals

### v2

@@ Appendix D Machine learning with the quantum reservoir
We then computed the mean squared error between the resulting $y_{\text{prediction}}$ and the label for the training point that produced the underlying reservoir feature $x$. Finally, we used back-propagation to compute the gradient for our linear layer. The linear layer was then updated using the ADAM optimizer [ADAMpaper] with the default settings of $\beta_{1}=0.9,\beta_{2}=0.999$ and a learning rate of $0.01$. For our reservoirs, we tried both methods of training the linear layer and used whichever yielded the best accuracies. Empirically, we found that while pseudo-inverse training was better in some cases, training the linear layer with back-propagation often yielded quite large accuracy advantages over pseudo-inverse.
Appendix E Supplementary information machine learning tasks
@@
In this section, we discuss about the algorithm for generating the dataset for the classification of digital modulation schemes on radio signals. The digital modulation scheme involves encoding sequences of binary values into the amplitude and phase of a radio signal for a fixed duration. The number of binary values encoded depends on the modulation scheme. For example, for BPSK (binary phase shift key), each symbol (change in property of the signal) encodes one bit of information. For $32$QAM (quadrature amplitude key), there are $32$ possible values, which allows each symbol to contain $5$ bits of information. For this task, we keep the symbol rate fixed across all the tasks. Moreover, the pulses generated by the arbitrary waveform generator (AWG) all occur at the baseband frequency. This signal is then upconverted to the frequency of the cavity before sent to the device. To generate the set of possible sequences, we randomly select each symbol with equal probability. This corresponds to the case of each possible binary string of digits encoded to be equally likely. Due to memory constraints on the AWG, we cannot output a continuous encoded signal for long durations, corresponding to the regime of large samples of the reservoir. We circumvent this constraint by realizing that, for this task, there are no correlations in the encoded binary digit sequence (since each symbol is equally likely). Therefore, the probability of a long binary digit sequence can be correctly emulated by sampling multiple short binary digit sequences and concatenating them together. For this task, we can simply achieve this by generating a signal with eight symbols, which is the number of symbols enter our QRC before its state is completely reset.
@@
Figure S14: Histogram of integrated value of the input signal for the noise classification task. In this work, we enforced a normalization condition on the amplitude of the filter functions. We set the normalization such that the long time integral of the signal corresponds to a value with zero mean and the same standard deviation. This is visually seen from the probability density function from the dataset of the classes of noisy signals. We do this to ensure that any reservoir which simply integrated the signal, before applying a non linear kernel, cannot classify the different signals. The fact that our reservoir is able to solve this task can therefore be associated with the continuous-time processing by the cross-Kerr interaction between the qubit and the cavity. Enforcing this normalization is mathematically equivalent to setting the DC component of the filter functions to be the same value in frequency space.
To generate the dataset describing the task of classifying noisy signals using the QRC (see Fig. 4), we start with emulating white noise. At each time step of the sampling rate of the AWG, we choose a value for in-phase and quadrature signals uniformly between the unit interval (up to an overall normalization). While this is limited to the sampling rate of the AWG (around $2\times 10^{9}$ samples per second), this is much larger than any relevant time scale of the experiment. Therefore the approximation of broadband white noise is appropriate to describe the effect of the signal on the system. We then apply “kernels” as a convolution in the time domain to each need seed of the white noise generated signal. This can also be thought of a bandpass filtering function in frequency domain. The classification task is then to identify the kernel. Each kernel is defined by a time domain function. The only hyper-parameter to describe each kernel is the overall scaling value. In this work, we set the DC component of this kernel in frequency domain to be the same for all classes (set to unit value without loss of generality). In the time domain, this corresponds to scaling the amplitude such that the area enclosed by the filter function in time is the same for all functions.. We do this to make sure that a direct integration of the signal over a time domain much longer than the correlation length introduced by the kernel, cannot distinguish the signals from each other (see Fig. S14). The above normalization ensures the random variable associated with this integrated value is the same for all distributions. Therefore, this ensures that any ability of the reservoir to classify the signals arises intrinsically from its computational capacity to distinguish short-time correlations (in this work we choose a correlation time scale of $50$ns and $600$ns, with kernel functions of Gaussian, Lorentzian, and the Inverse function: generating a total of $6$ classes.).
Appendix F Simulation of the quantum reservoir
@@ Introduction
Figure S15: Classification accuracies obtained from numerical simulations of the QRC. Numerical simulations of the quantum reservoir can help guide the expected performance of the system in experiment. Here, we simulate the three main tasks considered in the paper: Spiral (a time-independent input with the goal of classifying the arms of the spiral), Digital Modulation (a slow-varying, time-dependent signal with the goal of identifying the digital modulation scheme used in the signal), Correlated Noise (a fast-varying, time-dependent signal with the goal of classifying the “kernel” of the correlation function).
@@
To make the simulations efficient, we make certain assumptions on the quantum system. Firstly, we treat the reservoir controls of qubit rotations and conditional displacements with a “gate”-based unitary. However, to take into account the analog, continuous dynamical evolution implemented by the cross-Kerr interaction term in the Hamiltonian, the interval of the input into the system is implemented with the full time-dependent Hamiltonian evolution (using QuTiP’s “mesolve” functions). Finally, another approximation we make (in favor of simulation speed) is ignoring decoherence effects. To ensure this approximation is valid, we performed simulation with the stochastic wavefunction approach with photon loss and qubit dephasing rates measured in experiments [PhysRevLett.68.580]. We obtain differences in expected classification accuracies within error bars (which are obtained from different datasets from repeated simulations). This also gives us confidence that the role of decoherence in the system plays a minimal role in the computational capacity of the reservoir. The results of the simulations with the third central moment are plotted in Fig. S15. Interestingly, the performance as a function of the number of samples agrees to experiment within the same order of magnitude. This gives a good estimate for the experimental time required to produce a classification accuracy versus shots curves in experiments. For all three tasks, the reservoir approaches $100\%$ accuracy with sufficient samples or integration time of the input.
@@
Figure S16: Simulations comparing the performance of a continuous-time quantum reservoir (our work), and a discrete-time qubit-based quantum reservoir, based on a recently implemented protocol [7]. (a.) Schematic reservoir protocol for our experiment (on the left), versus that introduced in  [7], involving two qubits as the data qubits, and two as the ancillas. The reservoir consists of single qubit rotations interleaved with CNOT gates between the data qubits and data and ancilla qubits. In this simulation, we simulate the performance of the two reservoirs as a function of the delay $\tau$ between two durations of the input. Such a delay can be introduced by the finite pulse durations and latencies introduced by the FPGA. (b.) Classification accuracy curves for the two reservoirs as a function of shots, for different values of delays. The continuous time analog reservoir is much more robust to delays between inputs compared to the discrete time reservoir implementation. (c.) Plot of accuracies at $5000$ shots for the two reservoir implementations as a function of delay time. Experimentally relevant times include $4\mu s$ for our experiment, and around $8\mu s$ [hua2023exploiting] for the experimental realization of [7] on an IBM quantum computer.
@@
In this section, we benchmark the performance of our continuous-time-continuous variable QRC in comparison with other hardware implementations of reservoirs. To highlight the benefit of our QRC in processing time varying input signals, we compare the simulation of our reservoir with that of a recent QRC scheme involving repeated measurements on a multi-qubit based superconducting circuit quantum system [7]. For this comparison, we simulate the expected performance of both systems on the task of classifying different noise signals with the classes described in Fig. 4. While our reservoir can naturally interface with analog signal, this is not the case with the protocol introduced in [7]. For this simulation, the signal is sampled at discrete times and input to the system as a scalar parameter (one for the in-phase value and one for quadrature value). To highlight the advantage of our QRC, we slightly modify the task introduced in the Fig. 4. Here, we normalize the six filter functions such that the integral of the filter function in frequency domain is kept constant. We do this such that the standard deviation associated with the distribution of the sampled signal is the same across all signals. The only information distinguishing the signals is in the correlation between two close samples in time. To elucidate this reasoning, we simulate the performance of the two reservoirs as a function of the time duration in between two samples of the signal (in the case of the discrete qubit based reservoir) and integration windows (for our analog reservoir). Such a finite duration can arise from finite-pulse durations of reservoir protocols, qubit-measurement times, and the finite latency introduced by the classical FPGA processor. For example, for our experiment, this time is around $4\mu s$, mostly arising from the measurement of the qubit and the parity of the cavity. In experiment (Fig. 4), we had generated and timed the input wave forms such that the delay between inputs is essentially $0\mu s$. For a typical IBM quantum device with mid-circuit measurement, the protocol used in Ref. [7], the finite latency can be estimated to be around $8\mu s$ [hua2023exploiting]. The protocol for the discrete-time quantum reservoir is designed to only act on real-valued input signals. However, for a continuous signal in the rotating frame, we have both the in-phase and quadrature values. For the experimental quadrature, these values correspond to displacements on the oscillator in orthogonal directions. To extend the scheme presented in [7], we do the following minimal change: we interleave the between sample points of in-phase and quadrature values. We could have chosen these points with the delay of $\tau$ in between each. However, this might have had the effect of introducing twice the delay compared to the continuous time reservoir. Therefore, we chose the relaxed constraint of the input such that both the in-phase and quadrature values are chosen at the same point, with just a delay in between two different in-phase and quadrature points.
@@
Figure S17: Simulations of the performance of the cavity-and-qubit reservoir system with its components: just a cavity reservoir and just a qubit reservoir. (a.) The circuit diagrams of the three quantum reservoir simulated here. 1) The circuit diagram of the experimental QRC in this work: a cavity coupled to a qubit. The input is interfaced to the cavity, and the state of the qubit is measured using the standard dispersive readout technique with a readout resonator. 2) The circuit diagram of a cavity as a reservoir. In this case, the natural output of the cavity is the transmitted signal. 3) The circuit diagram of a qubit reservoir. In this case the input is coupled instead to the qubit. (b.) Simulation of the classification accuracy of the three different reservoirs for the three tasks considered in the paper. As a fair comparison, the output feature vector dimension is kept constant for all three reservoir. The experimental setup drastically outperforms either of its components, just a cavity (in (c.) and just a qubit (in (d.)), which highlights the important role of entanglement in classification accuracy.
@@
A cavity coupled to a qubit is a hardware efficient quantum system to perform reservoir computing on analog signals. In this section, we motivate this by simulating the performance of other natural choices of quantum reservoirs: a single qubit, and a single cavity. The protocols for these systems are inspired by what one can naturally perform in experiment. To make a reservoir with a cavity, we couple the input into the cavity (as is the case for the experimental design). To readout the cavity, we perform a transmission style Homodyne measurement, which infers the mean field value of the cavity. This is a continuous form of measurement, where the output feature is a time dependent radio frequency signal at the frequency of the cavity mode. Since the cavity is always in a coherent state, the output time trace is linearly dependent on the incoming signal. For a fair comparison, we only use a handful of values from the time trace (as many as the number of measurements in the experiment). While this might seem restrictive, we process this via the same method as the case of the experimental reservoir, by computing the functional definition of the central moments. This does not necessarily make sense for this protocol, since the outputs do not correspond to samples from a discrete probability distribution, but can nevertheless introduce non-linearities in the representation of the feature vector. These non-linearities can improve the performance of the reservoir beyond a linear layer. This is observed for the case of time-independent Spiral classification, where the cavity reservoir performs better than random. This performance is solely due to the “post-processing” of the output of the reservoir we adopt for our experiment. However, for time dependent tasks, the performance is hardly better than random.
@@
Another natural candidate is a single qubit reservoir. For this case, we directly interface the signal to the qubit. A qubit is able to naturally represent non-linear functions of the input, which can be intuitively seen by visualizing the action of a qubit rotation on a Bloch sphere. As a fair comparison, we choose the same qubit reservoir controls in experiment, which involve qubit pulses before, during and after the continuous input. The output is a string of binary outcomes of qubit measurements, which can be done experimentally with the use of a readout resonator. Each reservoir of the qubit lasts twice as long as the experimental QRC to obtain the same feature vector size. This is then processed the same way as the cavity-qubit coupled reservoir, before applying a trained linear layer. Interestingly, the qubit fails to perform better than random for the Spiral task. On the other hand, it is able to reach near $100\%$ accuracy for the time dependent signal classification tasks. The ability for even a single qubit to successfully perform a many-class classification task is illuminating at the remarkable processing capabilities of reservoir. However the total input signal required can be more than order of magnitude longer compared to that for the experimental QRC to achieve the same accuracy. The ability of a cavity coupled to a qubit system to perform significantly better than either of its components provides a clear picture of the important role entanglement can play.
@@
Figure S18: Exploring the computational capacity of more complex reservoirs, involving a single cavity mode coupled to multiple qubits. (a.) The reservoir protocol, before measurement. The protocol is inspired by naturally extending the protocol of the experiment. The state of each qubit, along with the parity of the cavity, is sampled afterwards. (b.) Schematic illustration of the state of the cavity after the generalized qubits conditioned cavity displacement implemented in this reservoir, for the case of two qubits. Here the cavity is displaced by a unique value of each $2^{N}$ combinations of qubit possibilities, for a reservoir with N qubits. (c.) To illustrate the computational capacity of such reservoir, we simulate the quantum system to obtain the classification accuracy as a function of the duration of signal received for the task of classifying correlated noise signals. Increasing the number of qubits generally increases the performance of the reservoir. For different values of shots of the reservoir, which corresponds a total time of input, we plot the classification error as a function of qubits. The case of zero qubits in the reservoir corresponds to just a cavity reservoir.
@@
Quantum reservoir computing is a promising paradigm in the NISQ era. It is therefore interesting to consider the potential benefits in performance with larger devices, which are within reach of today’s experimental capabilities. As a natural extension of our quantum reservoir, we consider a scenario of one continuous variable cavity mode dispersively coupled to multiple qubits. For simplicity, we assume the dispersive strength of each qubit to the cavity is the same. To motivate the capacity of such a reservoir, we simulate the system for up to four qubits to estimate the classification accuracy for the task of identifying correlated noise signals. The unitary protocol is illustrated in Fig S18 (a.). The protocol begins with a $\pi/2$ pulse on each qubit, which brings the state of the qubit onto the equator of the Bloch sphere. To entangle the qubits with cavity, we simulate the action of a generalized, multi-qubit-conditioned cavity displacement. This involves a displacement on the cavity, whose value is different for each of $2^{N}$ N-qubit possibilities. An example of the action of this operator is depicted in Fig S18 (b.) for the case of two qubits. Here, there are four possible qubit states, and each is associated with a displacement value of the four corners of the square. The displacement values were chosen somewhat arbitrarily, but serve to illustrate an efficient multi-component entanglement. For the case of two qubits, the real and imaginary components of the displacement where either ±$\pm 0.5$. For the case of three qubits, there are eight total possible states. The set of displacements chosen form a three-by-three grid state, ranging between ±$\pm 1$, excluding the center of this grid (which is centered at the origin). For the case of four qubits, a four-by-four grid uniformly distributed between ±$\pm 1.5$ covers all sixteen possibilities. The correspondence between qubit states and displacements was somewhat arbitrary - the motivation is that even without much design choice, a reservoir can successfully implement machine learning! For these simulations, each position of the grid is associated with a decimal value, increasing sequentially from left to right, starting from the top left and progressing towards the bottom right (starting with zero). This decimal value is the decimal representation of the qubit-state bit string that the displacement is conditioned on.
@@
Fig S18 (c.) is the classification accuracy as a function of the total time of input signal received for the reservoirs with different number of qubits, for the task of noise classification. While they all achieve essentially $100\%$ accuracy, the total time required to achieve this accuracy drops significantly with increasing number of qubits. Other than the single qubit reservoir (the experimental protocol), the performance of the reservoir is similar both at the low and high signal duration regime, differing only in the intermediate regime. The reason for the difference in behavior of the performance for the case of cavity coupled to a single qubit is the slight change in reservoir protocol. To accurately account for the experimental protocol, the state of the qubit is determined by the outcome of the parity measurement of the cavity. This was not implemented in the simulations for multiple qubits. This ends up improving the performance of this reservoir for this task in the low signal duration regime. However, in the higher signal duration regime, increasing the number of qubits increases the accuracy.
@@
The classification error as a function of number of qubits in the reservoir is plotted in Fig S18 (d.), including the case for just a cavity (zero number of qubits in the reservoir), for a select number of total shots of the entire reservoir. Very crudely, the error in classification seems to reduce exponentially with every additional qubit in the reservoir.
Appendix G Theoretical analysis of the expressivity of our QRC for time-independent signals